Constants Part A A charged paint is spread in a very thin uniform layer over the surface of a plastic sphere of diameter 14.0 cm, giving it a charge of 10.0 μC Find the electric field just inside the paint layer. N/C Submiteatses
Constants PartB ead in a very thin uniform layer over the here of diameter 14.0 cm, giving it a Find the electric field just outside the paint layer. N/C Submit
media%2Ffbf%2Ffbf97eaf-e435-40d6-8d60-7c
Answer #1

Part A

E = 0 as charge enclosed is zero.

Part B

E = kq / r2

here r = 14/2 = 7 cm = 0.07 m

E = 9e9 * 10e-6 / 0.072

E = 1.837e7 N/C

PART C

Here r = 0.07 + 0.095 = 0.165 m

E = 9e9 * 10e-6 / 0.1652

E = 3.305e6 N/C